# Adaptive Attention Distillation for Robust Few-Shot Segmentation under Environmental Perturbations

### v1

@@ I Introduction
Image segmentation, a fundamental task in computer vision, aims to precisely delineate object boundaries and plays a critical role in domains such as medical imaging and aerospace applications [8, 4, 46, 6, 15]. However, building large-scale training datasets remains expensive and time-consuming, as it requires extensive manual annotation of segmentation masks. This challenge has motivated research into few-shot segmentation (FSS), which learns to segment new object categories from only a handful of labeled examples.
A common solution is the pretrain–finetune paradigm, where pre-trained segmentation models (e.g., Swin Transformer [27], SAM [20]) are adapted to new tasks using limited samples. Yet under severe data scarcity, fine-tuning often leads to overfitting. FSS addresses this limitation by learning transferable knowledge from a small set of support images to segment unseen query images. Most FSS methods adopt a meta-learning framework with Siamese or prototypical architectures [39, 44, 32, 49, 25, 2, 7, 13], enabling the model to extract class-level representations shared across instances. Recent studies [48, 39] demonstrate that such methods achieve nearly $70$ mIoU in the 1-shot setting on general benchmarks including PASCAL [37] and COCO [33].
@@
To address these challenges, the Environment-robust Few-shot Segmentation (ER-FSS) task (see Fig. 1) is introduced to improve the resilience of FSS models under environmental perturbations. The task targets typical hard cases in query images arising from complex real-world conditions, such as motion blur, small objects, camouflaged targets, and occlusion of key features. To better mirror practical usage, images exhibiting these challenges serve as query images, while simpler, cleaner samples captured in controlled settings are used as support images. Based on this setup, the ER-FSS benchmark is constructed, covering six scenario types and eight datasets. Unlike conventional datasets, ER-FSS more faithfully reflects model performance, generalization, and robustness across diverse environmental variations and domains, offering a realistic benchmark for evaluating both general and FSS models.
On the ER-FSS benchmark, evaluations of state-of-the-art pretrain–finetune and FSS models reveal that their robustness under environmental variation remains far below practical requirements. Further analysis shows that perturbations amplify feature discrepancies between same-class targets in query and support images, leading to attention drift—the model fails to focus on the correct category or its key visual cues. To mitigate this issue, Adaptive Attention Distillation (AAD) is proposed. The method repeatedly contrasts semantic information between support and query images and progressively distills class-specific attention through multiple refinement stages. By consolidating critical semantic cues, AAD enhances target localization and segmentation accuracy in complex environments. Experimental results demonstrate that AAD consistently outperforms existing pre-trained and FSS models, achieving an average IoU improvement of 3.3%–8.75% over current state-of-the-art (SOTA) approaches.
@@ II-A Few-Shot Segmentation (FSS)
FSS [37, 33, 24, 32, 2, 7, 42, 49, 14, 26, 31] is characterized by the absence of target domain data during training. The goal is to segment query images from an unseen domain using only a few annotated support images. OSLSM [37] was the first to tackle this problem, computing classifier weights for each query-support pair at evaluation. Inspired by ProtoNet [41], most modern FSS methods follow a meta-learning framework with dual branches: one extracts prototypes from support images, the other processes query images. These approaches generally fall into two categories: prototypical feature learning and relation-based methods.
Prototypical feature learning improves prototype representations to better separate foreground and background, enabling more accurate similarity measurement between support and query images. For example, PANet [44] employs prototype learning and a non-parametric decoder to create a consistent metric space. Relation-based methods, on the other hand, focus on improved measures of similarity after feature extraction.
However, most methods overlook a practical challenge: query images are often more complex than support images. When query images lack clear category cues—such as when key attributes are heavily occluded—FSS models may fail to identify the target or incorrectly segment unrelated objects. This paper extends FSS to a multi-domain English context, aiming to improve model robustness across diverse real-world settings. To address this, this paper introduces a new benchmark dataset and propose a novel approach tailored for robust multi-domain FSS.
@@ II-B Related Datasets
To evaluate FSS model performance, most studies use PASCAL [37] and COCO [33], which cover common categories such as people, animals, vehicles, and indoor scenes, with COCO also including food, toys, and more. During testing, images for each novel class are randomly selected as support and query sets. However, these datasets do not adequately reflect the complexities of real-world applications.
To better assess FSS model generalization, the Cross-Domain Few-Shot Segmentation (CD-FSS) benchmark [21] was introduced, spanning the general domain (FSS-1000 [24]), medical (Chest X-ray [5]; ISIC [9]), and agricultural (Deepglobe [10]) domains. Despite this, CD-FSS remains limited in scope. Except for Deepglobe, FSS methods perform similarly on other datasets as in the general domain, indicating that CD-FSS does not fully capture the real challenges faced by current segmentation models. Like PASCAL and COCO, it also overlooks the impact of many difficult real-world cases on model generalization.
To address these gaps, we introduce the Environment-Robust Few-Shot Segmentation (ER-FSS) benchmark, designed to assess FSS under eight challenging scenarios, including camouflaged objects and small targets. ER-FSS provides a more comprehensive and realistic evaluation platform for FSS and broader segmentation models.
@@ III-A Overview of the Benchmark Dataset
In this work, we introduce an Environment-Robust FSS Benchmark (ER-FSS) benchmark dataset, designed to serve as a comprehensive evaluation platform for FSS models and a wide range of segmentation algorithms under diverse and realistic environmental conditions. Compared with previous benchmarks, ER-FSS offers several key advantages: (1) it covers a broader array of specialized domains; (2) it features more meticulous data cleaning and manual annotation; and (3) it explicitly distinguishes between easy and difficult samples according to environmental complexity—challenging real-world samples are designated as “query“ (evaluation) images, while simple or laboratory-scenario images are categorized as “support“ (known) images.
Specifically, for (1), the ER-FSS benchmark includes eight evaluation datasets spanning six major domains: Animals (biology, segmenting 18 animal categories), Lunar Terrain (astronomy, segmenting surface elevations and depressions), Polyps (medicine, segmenting colon polyps in colonoscopy images), Eyeballs (medicine, segmenting retinal blood vessels), Road Cracks (industry, segmenting cracks on road surfaces), Steel Defects (industry, segmenting surface defects on steel structures), Leaf Diseases (agriculture, segmenting diseased regions on leaves), and City Satellite (geography, segmenting nine categories of objects, including buildings and roads, in satellite imagery).
For (2), every image in the dataset has undergone manual inspection for quality and label accuracy, accompanied by rigorous data cleaning to ensure the highest annotation standards. For (3), we identified common real-world challenges such as high similarity between the target and background in color or shape, targets appearing too small due to long-distance imaging, elongated target shapes, image blurring, and occlusion or absence of key target attributes. We summarize these difficulties into five challenging characteristics: camouflaged objects, small targets, elongated targets, missing attributes, and image blurring. Accordingly, every image was manually annotated as either an “easy sample“ or a “hard sample“ and assigned to the “support set“ or “query set“, respectively.
@@ III-B Construction Process
Data Distribution and Statistics. To ensure the effectiveness of testing, the number of support images for all categories exceeds $20$, and the number of query images exceeds $10$. Moreover, to enhance dataset diversity, each dataset’s query images incorporate more than two challenging attributes. Additionally, we showcase visualization results of selected images from the Animal dataset after ViT [11] feature extraction with tSNE. As shown in Fig. 2, even within the same category—such as “snake,” “frog,” and “lizard”—the distribution of hard samples differs markedly from that of simple samples. Notably, hard samples of “frogs” are distributed more similarly to simple samples of “lizards” than to simple samples of their own class. This highlights a key challenge in real-world settings, where FSS and segmentation models often fail to identify targets correctly, or mistakenly classify them as background, resulting in segmentation errors. To address this, the ER-FSS benchmark proposed in this paper aims to offer a more realistic and application-oriented evaluation platform that better captures the complexities of real-world scenarios.
@@ IV-A Problem Setting
The environment-robust few-shot segmentation (ER-FSS) problem in this paper is formulated based on the classic FSS task. We assume a pre-training dataset $(X_{P},Y_{P})$ and a target domain dataset $(X_{T},Y_{T})$, where $X$ represents the input data distribution and $Y$ represents the corresponding segmentation labels. The model learns segmentation meta-knowledge from the extensive pre-training data $D_{P}\subseteq(X_{P},Y_{P})$ and is evaluated on the target data $D_{T}\subseteq(X_{T},Y_{T})$. Similar to FSS, in our ER-FSS setting, the input data distribution of the pre-training domain, $X_{P}$, differs from that of the target domain, $X_{T}$, and their label spaces are disjoint, i.e., $X_{P}\neq X_{T}$ and $Y_{P}\cap Y_{T}=\emptyset$.
The training set, denoted as $D_{\text{train}}$, is constructed from $(X_{P},Y_{P})$, while the testing set, $D_{\text{test}}$, is derived from $(X_{T},Y_{T})$. During the testing phase, we adhere to the episodic paradigm [43]. Specifically, for a given $N$-way $K$-shot learning task, $D_{\text{test}}$ comprises multiple episodes. Each episode is constructed with (1) a support set of simple samples, $S=\{({I}^{s}_{i},{M}^{s}_{i})\}_{i=1}^{N\times K}$, and (2) a query set of hard samples, $Q=\{({I}^{q}_{j},{M}^{q}_{j})\}_{j=1}^{n}$, where $I$ is an image, $M$ is the corresponding segmentation mask, and $n$ is the number of query images. Note that in our experiments, the training process for both the FSS baselines and our framework also follows the episodic paradigm. In contrast, when assessing pre-trained models based on transfer learning, the training phase adheres to their native single-branch, end-to-end approach.
@@ IV-B Method Overview
Current methods in FSS aim to extract knowledge of a new class from a support image to then segment a query image. However, these methods are challenged by the significant domain gap that arises when the target’s environment is complex, causing large intra-class variance between the support and query sets. For example, as illustrated in Fig. 2, a “rabbit” in a support image may be clearly depicted, while in a natural query scene, its features can be obscured by camouflage or occlusion. This variance leads to biased and non-generalizable knowledge being learned from the support image, causing models to fail at accurately localizing and segmenting difficult targets in query images.
To overcome this limitation, we introduce Adaptive Attention Distillation (AAD), a strategy designed to learn more robust class representations. AAD iteratively compares the core features of support and query images to “distill” a stable, generalizable class-level attention. This distilled attention effectively guides the model to focus on the target while differentiating it from the background. The adaptive nature of this process enables generalization to diverse and unseen environments, thus enhancing the overall environmental robustness of the FSS model. The proposed AAD framework, depicted in Fig. 3, is composed of four modules: a shared encoder, a correlation learner, an Adaptive Attention Distillation learner, and a decoder.
@@ IV-D Correlation Learner
As shown in Fig. 3(b), the correlation learning module is designed to establish an initial correspondence between the support and query images. This module is composed of multiple Aggregation Mask Modules (AMMs), each operating at a specific feature scale. The primary function of the AMM is to generate a coarse segmentation mask for the query image by transferring knowledge from the support set. This is achieved by leveraging a cross-attention mechanism, which has proven effective at identifying feature similarities across different inputs. The process begins with the AMM at each scale $i$ taking the support features $f^{s}_{i}$ and query features $f^{q}_{i}$ as input. The cross-attention mechanism computes the similarity between them, effectively assigning higher weights to regions in the query features $f^{q}_{i}$ that closely correspond to the target features present in $f^{s}_{i}$. Subsequently, the module incorporates the corresponding support mask $M^{s}_{i}$ to filter and refine these weighted features, producing a coarse mask that approximates the target’s location in the query image.
These multi-scale coarse segmentation maps serve a dual purpose. First, they are forwarded to the decoder and progressively upsampled to contribute to the final, high-resolution segmentation result. Second, and more critically for our AAD framework, they provide the initial target localization information required by the subsequent Adaptive Attention Distillation learner, which will further refine this understanding.
@@
In this formulation, the query features $f^{q}_{i}$ function as the ”Query” (Q), while the support features $f^{s}_{i}$ serve as the ”Key” (K). The dot product $f^{q}_{i}(f^{s}_{i})^{T}$ calculates a raw similarity matrix between every feature vector in the query and every feature vector in the support. By applying the softmax function, we obtain a set of attention weights where query features that are highly similar to support features receive larger values. This process effectively highlights regions in the query image that share semantic content with the support image.
The resulting attention map is then multiplied by the downsampled support mask $M^{s}_{i}$, which acts as the ”Value” (V). Since $M^{s}_{i}$ contains binary values indicating the target object’s location (1 for the target, 0 for the background), this final matrix multiplication effectively ”filters” the attention scores. It retains and aggregates the attention weights corresponding only to the target category, thereby producing a coarse probability map for the target’s location within the query image, which we denote as the coarse query mask $M^{q^{\prime}}_{i}$.
@@ IV-E Adaptive Attention Distillation Learner
While the Correlation Learner provides an initial, coarse localization of the target, its reliance on direct feature matching makes it susceptible to failure in ”hard” cases where significant appearance gaps exist between the support and query images (e.g., due to camouflage, motion blur, or viewpoint changes). To overcome this, a more robust mechanism is needed to distill the essential, class-discriminative semantics of the target category, independent of low-level feature variations. For this purpose, we introduce the Adaptive Attention Distillation (AAD) learner, as in Fig. 3(c). This module is designed to generate a set of compact, highly informative ”class queries” that encapsulate the core characteristics of the target category. Instead of performing dense, pixel-to-pixel comparisons, the AAD learner abstracts class-level information, enabling it to focus on the target’s fundamental properties while ignoring distracting background clutter and appearance shifts. The module consists of multiple Adaptive Query Generators (AQGs), which leverage a small set of learnable parameters to interact with and distill information from the foreground features of both the support and query images.
The process begins by isolating the foreground features, which is critical for ensuring the learner focuses exclusively on the target object. Using $f^{s}_{i}$ and $f^{q}_{i}$ from the encoder and the masks $M^{s}_{i}$ and $M^{q^{\prime}}_{i}$ (from the support set and the previous stage, respectively), we compute the foreground features $F^{s}_{i}$ and $F^{q}_{i}$ via element-wise multiplication:
@@
This operation effectively masks out the background, minimizing its influence on the subsequent learning process.
Next, we initialize a small set of $N$ learnable class queries, denoted as $q\in\mathbb{R}^{N\times l}$, where $l$ is a user-defined hidden dimension. In our experiments, we found that a small number of queries (e.g., $N<100$) is sufficient, making this a computationally lightweight operation. These queries act as ”information collectors,” tasked with distilling the most salient class semantics.
The core of the AQG is a cross-attention mechanism, but its application here is novel. The class queries $q$ act as the ”Query” (Q), while the support foreground features $F^{s}_{i}$ serve as the ”Key” (K), and the query foreground features $F^{q}_{i}$ serve as the ”Value” (V). This formulation is designed to answer the question: ”Which parts of the query foreground $F^{q}_{i}$ are most relevant, given the class context provided by the support foreground $F^{s}_{i}$?” The updated queries are computed as follows:
@@
where the output is passed through a LayerNorm and a simple MLP (consisting of two linear layers and a ReLU activation) to refine the representation and reduce its dimensionality back to the original $N\times l$.
It is crucial to note that this process is fundamentally different from dense feature matching. The lightweight queries $q$ do not learn a pixel-wise correspondence. Instead, they treat the entire set of support foreground features $F^{s}_{i}$ as a unified representation of the target class. This abstraction is key to the method’s robustness; it allows the queries to capture the essential what (the class identity) rather than the incidental where (the exact pixel locations), making the learned representation resilient to intra-class variations and noise in the coarse query mask $M^{q^{\prime}}_{i}$.
Furthermore, we leverage the hierarchical nature of features extracted by deep networks. Low-level features (from earlier layers, e.g., at scale 1/8) typically encode rich edge and texture details, while high-level features (from deeper layers, e.g., at scale 1/32) capture more abstract semantic information. To harness this, the query distillation process is performed iteratively across the different scales. The updated queries from one scale, $\tilde{q}_{i}$, become the input queries for the next scale. This allows the class queries to progressively refine their understanding, starting with general structural information and culminating in high-level semantic knowledge. The final output, $\hat{q}$, is a set of highly discriminative class queries that have distilled the core, environment-invariant essence of the target category, ready to guide the final segmentation decoder.

### v2

@@ I Introduction
Image segmentation serves as a cornerstone in computer vision, underpinning critical applications ranging from medical diagnosis to aerospace analysis [8, 4, 50, 6, 16]. However, the prohibitive cost of acquiring pixel-level annotations for large-scale datasets has catalyzed research into few-shot segmentation (FSS), which aims to segment novel categories given only a handful of labeled examples. This data-efficient paradigm has proven particularly valuable in specialized domains where labeled data is scarce. For instance, in medical imaging, recent methods [11, 44] leverage anatomical priors or volumetric consistency to handle diverse modalities; in remote sensing, approaches [54, 55] have been tailored to address significant scale variations and complex geospatial contexts; similarly, in industrial inspection, FSS [58] are increasingly adopted for detecting surface defects under limited sample conditions.
A prevalent strategy for label-efficient segmentation is the pretrain–finetune paradigm, wherein models pre-trained on large-scale data (e.g., Swin Transformer [28], SAM [20]) are adapted to downstream tasks. However, in regimes of severe data scarcity, fine-tuning is prone to overfitting. FSS circumvents this limitation by learning transferable patterns from a limited set of support images to segment unseen query instances. The majority of FSS approaches leverage a meta-learning framework underpinned by Siamese or prototypical architectures [42, 48, 33, 57, 25, 2, 7, 14]. Recent advancements [52, 42] have demonstrated the efficacy of these methods, achieving nearly $70\%$ mIoU in the 1-shot setting on standard benchmarks such as PASCAL-5i [40] and COCO-20i [34].
@@
To address these challenges, the Environment-robust Few-shot Segmentation (ER-FSS) task (see Fig. 1) is introduced to improve the resilience of FSS models under environmental perturbations. The task targets typical hard cases in query images arising from complex real-world conditions, such as motion blur, small objects, camouflaged targets, and occlusion of key features. To better mirror practical usage, images exhibiting these challenges serve as query images, while simpler, cleaner samples captured in controlled settings are used as support images. Based on this setup, the ER-FSS benchmark is constructed, covering six scenario types and eight datasets. Unlike conventional datasets, ER-FSS more faithfully reflects model performance, generalization, and robustness across diverse environmental variations and domains, offering a realistic benchmark for evaluating both pre-train segmentation and FSS models.
Evaluations on the ER-FSS benchmark demonstrate that both state-of-the-art (SOTA) pretrain–finetune and FSS models exhibit limited robustness under environmental variations, falling significantly short of practical deployment requirements. Further analysis reveals that such perturbations exacerbate feature discrepancies between intra-class targets in the query and support sets. This divergence precipitates attention drift, wherein the model fails to lock onto the correct category or misses critical visual cues.
To mitigate this issue, Adaptive Attention Distillation (AAD) is proposed. The method repeatedly contrasts semantic information between support and query images and progressively distills class-specific attention through multiple refinement stages. By consolidating critical semantic cues, AAD enhances target localization and segmentation accuracy in complex environments. Experimental results demonstrate that AAD consistently outperforms existing pre-trained and FSS models, achieving an average IoU improvement of 3.3%–8.75% over current SOTA approaches.
@@ II-A Few-Shot Segmentation (FSS)
Standard FSS [40, 34, 24, 33, 2, 7, 46, 57, 15, 27, 32] typically follows a meta-learning paradigm involving two distinct phases: training on source-domain base classes and inference on target-domain novel classes. During the training phase, the model learns generalizable feature representations from abundant annotated data of base classes. In the inference phase, the goal is to segment query images containing unseen categories, guided by only a few annotated support samples from the target domain.
OSLSM [40] pioneered this field by dynamically generating classifier weights for each query-support pair. Inspired by Prototypical Networks [45], most contemporary FSS methods adopt a dual-branch meta-learning framework, broadly categorized into prototypical feature learning and relation-based methods. The former, exemplified by PANet [48], focuses on optimizing prototype representations to better separate foreground from background, while the latter emphasizes refining similarity metrics between extracted features.
However, existing approaches often overlook a critical practical asymmetry: query images are frequently significantly more complex than support samples. When query targets suffer from severe occlusion or lack distinct category cues, standard models struggle to accurately localize the object. To address this, we extend FSS to a robust multi-domain setting, introducing a new benchmark and a novel approach specifically designed to enhance generalization across diverse real-world environments.
@@ II-B Related Datasets
Standard FSS evaluations predominantly rely on PASCAL-5i [40] and COCO-20i [34], which comprise common object categories. During testing, support and query sets are randomly sampled from novel classes; however, these datasets largely fail to capture the complexity inherent in real-world deployment. To address generalization, the Cross-Domain Few-Shot Segmentation (CD-FSS) benchmark [21] was proposed, incorporating diverse domains such as medical imaging [5, 9] and agriculture [10].
Yet, CD-FSS remains limited in scope. With the exception of Deepglobe, performance on CD-FSS datasets mirrors that of general domains, suggesting it does not impose sufficient difficulty to expose the vulnerabilities of modern models. Crucially, like its predecessors, CD-FSS overlooks severe environmental perturbations encountered in the wild. To bridge this gap, we introduce the Environment-Robust Few-Shot Segmentation (ER-FSS) benchmark. Spanning eight challenging scenarios—including camouflaged objects and small targets—ER-FSS provides a rigorous and comprehensive platform to evaluate the robustness of FSS models under realistic, adversarial conditions.
@@ III-A Overview of the Benchmark Dataset
In this work, we introduce the Environment-Robust FSS (ER-FSS) benchmark, a comprehensive platform designed to evaluate segmentation algorithms under diverse, realistic conditions. ER-FSS distinguishes itself from prior benchmarks through three key advantages: broad domain coverage, meticulous annotation quality, and a rigorous stratification of environmental complexity.
First, ER-FSS spans six distinct domains across eight datasets, covering: Biology (18 animal categories), Astronomy (lunar terrain features), Medicine (polyps and retinal vessels), Industry (road cracks and steel defects), Agriculture (leaf diseases), and Geography (satellite imagery of urban structures).
Second, to ensure high-quality supervision, every image underwent rigorous manual inspection and cleaning. Third, and most critically, ER-FSS explicitly models the domain gap inherent in real-world applications. We categorize images based on environmental difficulty: clear, laboratory-quality samples are designated as the support set, while challenging, in-the-wild samples serve as the query set. These challenges are formalized into five characteristics: camouflaged objects, small targets, elongated structures, missing attributes, and motion blur. This structured split ensures that models are evaluated on their ability to generalize from simple examples to complex, adversarial scenarios.
@@ III-B Construction Process
Data Distribution and Statistics. To ensure robust evaluation, we maintain a minimum of 20 support images and 10 query images per category, with query samples incorporating multiple challenging attributes to enhance diversity. We validate the proposed difficulty stratification via t-SNE visualization of ViT [12] features, as shown in Fig. 2. The analysis reveals a significant distribution shift between easy (support) and hard (query) samples within the same category. For instance, hard samples of ”frog” cluster closer to simple ”lizard” samples than to their own class prototypes. This intra-class divergence and inter-class confusion highlight the critical real-world challenge where models misclassify targets due to feature drift.
Furthermore, ER-FSS captures domain-specific imbalances inherent in practical applications. The challenge distribution varies naturally across datasets: industrial and agricultural subsets (e.g., Road Cracks, Steel Defects, Leaf Diseases) are dominated by slender structures and motion blur; biological data (Animals) emphasizes camouflaged and partially occluded targets; medical imagery (Eyeballs, Polyp) suffers from lighting variations and blur; while remote sensing domains (Lunar Terrain, City Satellite) predominantly feature small-scale objects. This heterogeneity ensures that ER-FSS provides a holistic assessment of model robustness across varying environmental complexities.
@@ IV-A Problem Setting
We formulate the Environment-Robust Few-Shot Segmentation (ER-FSS) problem building upon the classic FSS framework. Given a source domain $\mathcal{D}_{\text{train}}=\{(X_{P},Y_{P})\}$ and a target domain $\mathcal{D}_{\text{test}}=\{(X_{T},Y_{T})\}$, we assume a distributional shift between domains ($X_{P}\neq X_{T}$) and disjoint label spaces ($Y_{P}\cap Y_{T}=\emptyset$). The model acquires meta-knowledge on $\mathcal{D}_{\text{train}}$ and is evaluated on $\mathcal{D}_{\text{test}}$ following the episodic paradigm [47].
Specifically, for an $N$-way $K$-shot task, each episode comprises a support set $\mathcal{S}=\{(I^{s}_{i},M^{s}_{i})\}_{i=1}^{N\times K}$ and a query set $\mathcal{Q}=\{(I^{q}_{j},M^{q}_{j})\}_{j=1}^{n}$. A critical distinction in our ER-FSS setting is the stratification of difficulty: $\mathcal{S}$ is composed exclusively of simple samples , whereas $\mathcal{Q}$ consists of hard samples. This asymmetric design rigorously evaluates the model’s ability to generalize from ideal references to realistic, challenging environments.
Note that in our experiments, we adopt distinct training protocols tailored to each model architecture. For FSS baselines and our method, training strictly follows the episodic meta-learning paradigm. Conversely, for pre-trained models based on transfer learning (e.g., SAM), we adhere to their native single-branch, end-to-end approach. Specifically, these models leverage general representations learned from large-scale datasets (e.g., ImageNet). During the $N$-shot evaluation, we adapt them to the target task by fine-tuning the decoder on the provided $N$ support images, while keeping the backbone encoder frozen to preserve generalization.
@@ IV-B Method Overview
Current FSS methods typically transfer class knowledge from support images to segment query targets. However, performance degrades significantly when environmental complexity introduces large intra-class variance. For instance, as illustrated in Fig. 2, while a support “rabbit” may be distinct, its query counterpart is often obscured by camouflage or occlusion. This discrepancy leads to the extraction of biased prototypes, causing models to fail in accurately localizing targets within challenging scenarios.
To mitigate this, we introduce Adaptive Attention Distillation (AAD), a strategy designed to cultivate robust class representations. AAD iteratively aligns core features between support and query images to “distill” stable, class-level attention. This mechanism effectively suppresses background noise and highlights target regions, thereby enhancing generalization across diverse, unseen environments. As depicted in Fig. 3, our framework comprises four key modules: a shared encoder, a correlation learner, the AAD learner, and a decoder.
@@ IV-D Correlation Learner
As shown in Fig. 3(b), the correlation learning (CL) module is designed to establish an initial correspondence between the support and query images. This module is composed of multiple Aggregation Mask Modules (AMMs), each operating at a specific feature scale. The primary function of the AMM is to generate a coarse segmentation mask for the query image by transferring knowledge from the support set. This is achieved by leveraging a cross-attention mechanism, which has proven effective at identifying feature similarities across different inputs. The process begins with the AMM at each scale $i$ taking the support features $f^{s}_{i}$ and query features $f^{q}_{i}$ as input. The cross-attention mechanism computes the similarity between them, effectively assigning higher weights to regions in the query features $f^{q}_{i}$ that closely correspond to the target features present in $f^{s}_{i}$. Subsequently, the module incorporates the corresponding support mask $M^{s}_{i}$ to filter and refine these weighted features, producing a coarse mask that approximates the target’s location in the query image.
These multi-scale coarse segmentation maps serve a dual purpose. First, they are forwarded to the decoder and progressively upsampled to contribute to the final, high-resolution segmentation result. Second, and more critically for our AAD framework, they provide the initial target localization information required by the subsequent AAD learner, which will further refine this understanding.
@@
In this formulation, the query features $f^{q}_{i}$ function as the “Query” (Q), while the support features $f^{s}_{i}$ serve as the“Key” (K). The dot product $f^{q}_{i}(f^{s}_{i})^{T}$ calculates a raw similarity matrix between every feature vector in the query and every feature vector in the support. By applying the softmax function, we obtain a set of attention weights where query features that are highly similar to support features receive larger values. This process effectively highlights regions in the query image that share semantic content with the support image.
The resulting attention map is then multiplied by the downsampled support mask $M^{s}_{i}$, which acts as the “Value” (V). Since $M^{s}_{i}$ contains binary values indicating the target object’s location (1 for the target, 0 for the background), this final matrix multiplication effectively “filters” the attention scores. It retains and aggregates the attention weights corresponding only to the target category, thereby producing a coarse probability map for the target’s location within the query image, which we denote as the coarse query mask $M^{q^{\prime}}_{i}$.
@@ IV-E Adaptive Attention Distillation Learner
While the CL provides an initial, coarse localization of the target, its reliance on direct feature matching renders it susceptible to failure in “hard” cases characterized by significant appearance discrepancies (e.g., camouflage, motion blur, or viewpoint shifts). To transcend these limitations, a mechanism is required to distill essential, class-discriminative semantics independent of low-level variations. To this end, we introduce the AAD learner, illustrated in Fig. 3(c). This module generates a set of compact, highly informative “class queries” that encapsulate the core characteristics of the target category. Unlike dense pixel-to-pixel comparisons, the AAD learner abstracts class-level information, focusing on fundamental target properties while suppressing distracting background clutter. The module is composed of multiple Adaptive Query Generators (AQGs), which progressively interact with and distill information from the foreground features of both support and query images.
The process commences by isolating foreground features to ensure the learner focuses exclusively on the target object. Using feature maps $f^{s}_{i}$ and $f^{q}_{i}$ from the encoder, alongside masks $M^{s}_{i}$ and $M^{q^{\prime}}_{i}$ (derived from the support set and the preceding stage, respectively), we compute the foreground representations $F^{s}_{i}$ and $F^{q}_{i}$ via element-wise multiplication:
@@
This operation effectively masks out the background, minimizing its influence on the subsequent distillation process.
Distinct from traditional approaches that rely on static parameters, our query initialization is dynamic. We begin by generating a set of $N$ random vectors, denoted as $q_{in}\in\mathbb{R}^{N\times l}$, where $l$ is the hidden dimension. In our experiments, a small number of queries (e.g., $N=15$) proves sufficient. Although these queries originate from a random distribution, they function as adaptive “information collectors.” Through the forward pass, they act as “learnable” entities in the sense that they adaptively aggregate and refine semantic information from the support and query features, evolving from random noise into discriminative descriptors.
The core of the AQG is a cross-attention mechanism designed to extract relevant features. Here, the class queries $q$ serve as the “Query” (Q), the support foreground features $F^{s}_{i}$ act as the “Key” (K), and the query foreground features $F^{q}_{i}$ serve as the “Value” (V). This formulation addresses the question: “Which parts of the query foreground $F^{q}_{i}$ are most relevant, given the class context provided by the support foreground $F^{s}_{i}$?” The updated queries are computed as follows:
@@
where the output undergoes layer normalization and projection via an MLP to restore the dimensionality to $N\times l$. This design allows the queries to abstract the essential what (class identity) rather than the incidental where (pixel location), ensuring robustness against intra-class variations and noise in the coarse mask $M^{q^{\prime}}_{i}$.
Crucially, we exploit the hierarchical nature of deep features to progressively refine these queries. The process is iterative across different feature scales. The initially random query set $q_{in}$ first enters the AAD module at the $\frac{1}{8}$ scale, interacting with low-level texture and edge details to form the preliminary feature $q_{1}$. Subsequently, $q_{1}$ propagates to the second and third AAD modules, interacting further with features at the $\frac{1}{16}$ and $\frac{1}{32}$ scales, respectively. This multi-stage evolution allows the queries to integrate information ranging from structural details to abstract semantics. Ultimately, the final output $\hat{q}$ represents a set of highly focused, environment-invariant target attention features, ready to guide the final segmentation decoder.